solve the system of equations by graphing. 3x-2y=18, -3x-y=0
Answer by mananth(16949)

3x-2y=18-----------(1)
-3x-y=0-------------(2)
equation1
when x= 0 y= -9 ( 0 , -9 )
when y = 0 x= 6 ( 6 , 0 )
Equation 2
when x= 0 y= 0 ( 0 , 0 )
when y = 0 x= 0 ( 0 , 0 )
Plot these points and drwa the line graphs. The point of intersection of the lines is the solution of the system.
To draw the line y=-3x calculate the rise / run slope is -3.


The solution is (2,-6)
m.ananth@hotmail.ca